# hdMTD: An R Package for High-Dimensional Mixture Transition Distribution Models

### v1

@@ 1 Introduction
Categorical time series with long-range dependencies are ubiquitous in nature. Today’s weather depends not only on the weather of the previous day, but also on what the weather was like a year ago. In large language models, incorporating relevant words from distant positions in the input sequence can substantially enhance next-word prediction (more than $10^{5}$ for ChatGPT 4o). Economic indicators’ dynamics depends on events that happen at different temporal scales, including long-range dependencies. Modeling these phenomena using stochastic processes with possibly long-range dependencies is natural. Given the complex nature of some of these phenomena, often non-parametric modeling is desirable. Nevertheless, non-parametric modeling, like high-order Markov chains, generally requires estimating joint probability distributions, which imposes significant constraints on the size of the past that can be included in the model before the estimation becomes inaccurate. To ameliorate the problem, Raftery (1985) introduced the mixture transition distribution (MTD) model, a subclass of finite-order Markov chains, as a parsimonious non-parametric model for categorical time series. The MTD framework successfully reduces the number of parameters required for estimation, alleviating the complexity associated with high-order dependencies. Despite this advancement, accurately estimating time series dependencies of an MTD model that stretch far into the past remained limited. Recently, Ost and Takahashi (2023) introduced a computationally and statistically efficient solution to the problem of estimating time-series long-range dependencies if the time series is a realization of an MTD model in which the number of relevant lags in the past is not too big. This opened the possibility to make non-parametric inferences for the class of high-order MTD models. In this article, we introduce hdMTD, an R package allowing us to estimate parameters of such high-dimensional Markovian models.
@@
The core of the hdMTD package lies in the algorithms that, given a sample from an MTD chain, can retrieve the relevant set of past lags using two different methods: (1) the Bayesian information criterion (BIC) algorithm, which is a classical model selection method; (2) the forward stepwise and cut (FSC) algorithm, which is a computationally and statistically efficient algorithm for estimating the set of relevant lags even when the temporal distance between the lags is close to the sample size. Once the relevant lags are specified, hdMTD package can also compute the maximum likelihood estimate (MLE) of the transition probabilities and estimate high-order MTD parameters using the expectation-maximization (EM) algorithm. To allow for Monte Carlo experiments, it is necessary to sample from the invariant distribution of the process. For low-order Markov chains, traditionally we rely on running the process with fixed pasts and wait until the process approximate the invariant measure. When the order of the chain is large, this approach can be inefficient as it is difficult to decide whether the joint probability measure is close the to the invariant distribution. A more robust approach is to sample exactly from the invariant distribution. Algorithms that allow to draw samples from the invariant distribution of a process are called perfect sampling algorithms (see (Fernández et al., 2001) for a comprehensive introduction on the subject). Our package allows to simulate MTD chains of any order from its invariant distribution using the perfect sampling algorithm, which is novel to the best of our knowledge .
@@
A brief review of related computational packages helps to place hdMTD in context. Markovian models have a long history and broad applications, nevertheless, practical tools for inference, particularly for mixture transition distribution (MTD) models, remain relatively limited. Traditional statistical software (e.g., SPSS (IBM Corp., 2023), Stata (StataCorp, 2023)) do not provide native, dedicated support for Markov models, while SAS (SAS Institute Inc., 2023) offers limited functionality, primarily for homogeneous chains. Consequently, researchers increasingly rely on open-source environments like R (R Core Team, 2024) for advanced Markovian analysis. In R, the march (Maitre et al., 2020) (Berchtold et al., 2020) package provides dedicated tools for MTD model estimation, offering likelihood based optimization for categorical stochastic processes. This package has become a reference for MTD applications, supporting homogeneous Markov chains, hidden Markov models, and other related architectures. Its companion package GenMarkov (Vasconcelos and Damasio, 2023) extends this to multivariate MTDs. Notably, none of these tools provide perfect simulation algorithms of MTD processes given parameters. For general Markov chain analysis, the markovchain (Spedicato, 2017) package analyzes properties (e.g., stationarity) and experimentally fits $k$-step Markov chains, while DTMCPack (Nicholson, 2022) focuses on simulation from user-defined matrices. The depmixS4 (Visser and Speekenbrink, 2010) package offers tools for fitting standard and hidden Markov models with both continuous and categorical responses. Outside R, specialized tools like Stata’s markov module (Cox, 2002) implement basic discrete-time Markov chain analysis. More recently, developments in Julia (Bezanson et al., 2017) and Python (Python Software Foundation, 2023) have started to offer basic Markovian model support, but tools specific for MTDs remain largely unavailable.
@@ 2.3 Mixture transition distribution (MTD) models
An MTD (Raftery, 1985; Berchtold and Raftery, 2002) of order $d\in\mathbb{Z}^{+}$ is a Markov chain of order $d$ with state space $\mathcal{A}$, where the transition probabilities can be expressed as
@@
Henceforth, an index $j\in\operatorname{\llbracket}-d,0\operatorname{\rrbracket}$ will be called the j-th lag of the model. Throughout this article, for $j\in\operatorname{\llbracket}-d,-1\operatorname{\rrbracket}$, we denote by $p_{j}=\{p_{j}(a|b):\ a,b\in\mathcal{A}\}$, the $|\mathcal{A}|\times|\mathcal{A}|$ stochastic matrix representing the relationship between the $j$-th lag and the present. Equation (2) represents what Berchtold and Raftery (2002) called a multimatrix MTD model, since the stochastic matrices $p_{j}$ are allowed to be different from each other. Note that a multimatrix MTD model may have up to $d+(|\mathcal{A}|-1)(1+d|\mathcal{A}|)$ parameters. Hence, for such Markovian models the number of parameters grows linearly with the order $d$, unlike general Markov chains for which the number of parameters grows exponentially with the order. Note also that there are even more parsimonious MTD models, such as the single matrix MTD, where all lags influence the distribution of the present state equally (i.e., given any $a,b\in\mathcal{A}$, $p_{j}(a|b)=p(a|b)$ for all $j\in\operatorname{\llbracket}-d,-1\operatorname{\rrbracket}$). This latter MTD model has only $d+(|\mathcal{A}|-1)(1+|\mathcal{A}|)$ parameters.
@@ 3.2 Empirical probabilities
One can show that, when $\Lambda\subseteq S$, the empirical conditional probability $\hat{\operatorname{\mathsf{P}}}_{n}(a|x_{S})$ defined in (7) converges in probability to the conditional probability $\operatorname{\mathsf{P}}(a|x_{\Lambda})$ as the sample size $n\to\infty$, provided that any two vectors $x,y\in\mathcal{A}^{d}$ of $d$ past states communicate, meaning that the MTD model generating the sample is irreducible (see, e.g., Finesso (1991, Chapter 3)). Under this irreducibility assumption and if $\hat{\Lambda}_{n}=\hat{\Lambda}_{n}(\mathcal{S}_{n})$ is a consistent estimator of $\Lambda$ (i.e., the probability of the event $\{\hat{\Lambda}_{n}\neq\Lambda\}$ vanishes as the sample size $n\to\infty$), then we also have
@@ 3.3 BIC estimator
Model selection through the BIC (Schwarz, 1978) is a well-known method. The difference between BIC values of two models approximates twice the logarithm of the Bayes factor between them. The Bayes factor represents the evidence, provided by the data, in favor of one model against another (Kass and Raftery, 1995) and is known for its validity for comparison of multiple and non-nested models. Also, the BIC has been proven to be a consistent method for estimating the order of Markov chains (Katz, 1981; Csiszár and Shields, 2000).
@@ 3.4 CUT estimator
Ost and Takahashi (2023) proposed a consistent estimator for the set of relevant lags of an MTD model based on pairwise comparisons of empirical conditional probabilities corresponding to compatible pasts. In this article, we call it CUT estimator [^2] In Ost and Takahashi (2023), this estimator based on pairwise comparisons is called PCP estimator.. Given a subset $S\subseteq\operatorname{\llbracket}-d,-1\operatorname{\rrbracket}$ and a lag $j\in S$, two sequences $x_{S},y_{S}\in\mathcal{A}^{S}$ are called $(S\setminus\{j\})$-compatible if $x_{k}=y_{k}$ for all $k\in S\setminus\{j\}$. From our discussion at the end of Section 2.3, if $j\in\Lambda\subseteq S$ there must exist a pair of $(S\setminus\{j\})$-compatible pasts $x_{S},y_{S}\in\mathcal{A}^{S}$ such that the total variation distance between $\operatorname{\mathsf{P}}(\cdot|x_{S})$ and $\operatorname{\mathsf{P}}(\cdot|y_{S})$ is strictly positive. Hence, the CUT estimator runs across all $j\in S$, and tests if there is a pair of $(S\setminus\{j\})$-compatible pasts $x_{S},y_{S}\in\mathcal{A}^{S}$ for which
@@
with $\xi>0$, $\alpha>0$ and $\mu\in(0,3)$ such that $\mu>\psi(\mu)=e^{\mu}-\mu-1$. The threshold $t_{n}(x_{S},y_{S})$ has been derived through martingale concentration inequalities (see Ost and Takahashi (2023, Appendix B)). An interesting feature of this threshold is that it is adapted to each realization of sample $\mathcal{S}_{n}$ of the MTD model, i.e., it may change depending on the realization of the sample. Using uniform thresholds that do not depend on the sample typically lead to statistical procedures that either underestimate or overestimate the set of relevant lags.
The CUT estimator is computed in the function hdMTD_CUT. In Ost and Takahashi (2023), a few theoretical guarantees of the CUT estimator have been established. Let us highlight two of them. First, the CUT estimator can be implemented with $O(|\mathcal{A}|^{2}|S|(n-d))$ operations (see Ost and Takahashi (2023, Remark 2-(c))). Second, in Ost and Takahashi (2023, Corollary 4), the CUT estimator is shown to be consistent even when the order $d$ is large (e.g. $d=n\beta$ for some $\beta\in(0,1)$), provided that $\Lambda\subseteq S$ and $|S|$ is relatively small with respect to the sample size $n$ (i.e., $|S|=O(\log(n))$). However, this assumption is too restrictive and might be difficult to be justified in practice. To circumvent this issue, one can use the FS or FSC estimators described subsequently.
@@ 3.5 FS and FSC estimators
The FS estimator (Ost and Takahashi, 2023) is defined iteratively as follows. At each step, it receives a subset $S\subseteq\operatorname{\llbracket}-d,-1\operatorname{\rrbracket}$ as input and computes, based on the sample $\mathcal{S}_{n}$, the empirical estimate $\hat{\nu}_{n,j,S}$ of the quantity $\bar{\nu}_{j,S}=\mathbb{E}(\nu_{j,S}(X_{S})),$ for each lag $j\in S^{c}$. Here, for each $x_{S}\in\mathcal{A}^{S}$,
@@
As indicated in Ost and Takahashi (2023, Remark 8 - (c)), one can compute the FS estimator with $O(|A|^{3}\ell(n-d)(d-(\ell-1)/2)$ operations. Moreover, Theorem 2 of Ost and Takahashi (2023) states that, on a certain event and for a suitable chosen $\ell$, we have that $\Lambda\subseteq\hat{\Lambda}_{n}^{\rm FS}$. In hdMTD package, $\ell$ is a user input, and in practice, we verify that if $\ell\geq|\Lambda|$, then $\Lambda\subseteq\hat{\Lambda}_{n}^{\rm FS}$ with high probability. This suggests that the FS output might be a reasonable choice for input set $S$ of both BIC and CUT estimators. Let us mention that using the FS output as the CUT estimator input constitutes what Ost and Takahashi (2023) refer to as the FSC estimator. Roughly speaking, the FSC estimator is proved to be consistent even when the underlying MTD model has a large order $d$, as long as the size of the set of relevant lags $\Lambda$ grows at most logarithmically with the sample size $n$. As opposed to the CUT estimator, the FSC estimator does not require any prior knowledge of a sufficiently small subset $S$ containing the set of relevant lags $\Lambda$, which is a huge advantage in practice.
@@ 4 Algorithm to exactly simulate MTD models
There is no exact simulation algorithm implemented in any of the packages for MTD models that we are aware of, namely, the packages march and GenMarkov. In the hdMTD, the above exact simulation algorithm is implemented in the function perfectSample. Our algorithm is inspired by (Comets et al., 2002) - where the authors present an algorithm to exactly simulate processes with long memory. In that paper, the authors also discuss the differences and similarities of their algorithm with respect to other classical exact simulation algorithms such as coupling from the past algorithm introduced by (Propp and Wilson, 1996). We also refer to (Fernández et al., 2001) for a comprehensive introduction to perfect simulation algorithms.
@@ 5 Using hdMTD
R> install.packages("hdMTD") and loaded with:
@@ 5.1 Data generation
Suppose we want to obtain a sample from an MTD model with state space $\mathcal{A}=\{0,1\}$ where the set of relevant lags is $\Lambda=\{-30,-15,-1\}$. First, we need to specify such a model. We do so with the MTDmodel function. This function creates a class MTD object whose attributes are all the necessary parameters for defining an MTD model, and the resulting transition matrix P. The only inputs that are absolutely necessary are the set of relevant lags (Lambda) and the state space (A). Note that while $\Lambda\subset\mathbb{Z}^{-}$, the argument Lambda of the function MTDmodel must be in $\mathbb{Z}^{+}$. Hence, the input set Lambda has to be interperted as the set $\Lambda$. The user can also input the MTD weights (a number lam0 representing $\lambda_{0}$ and a vector lamj representing the values $\lambda_{j}$ for all $j\in\Lambda$), the past conditional distributions (a list pj where each entry is a matrix representing the matrices $p_{j}$, $j\in\Lambda$), and/or the independent distribution (a vector p0, representing $p_{0}(a)$, $a\in\mathcal{A}$). If not provided, these parameters are sampled uniformly and normalized [^3] For sampling $p_{0}(a)$, $a\in\mathcal{A}$, the function first sets p0<-runif(length(A)), and then renormalizes the vector $p_{0}$ to obtain a probability measure by defining p0<-p0/sum(p0). The same procedure is done for the weights (lam0 and lamj). For sampling $p_{j}(a|b)$, $a,b\in\mathcal{A}$, $j\in\Lambda$, the function sets a list called pj with $|\Lambda|$ matrices, where each row of each matrix is sampled in the same way as p0. .
R> set.seed(11)R> Lambda <- c(1, 15, 30)R> A <- c(0, 1)R> lam0 <- 0.01R> lamj <- c(0.39, 0.3, 0.3)R> p0 <- c(0.5, 0.5)R> MTD <- MTDmodel(Lambda = Lambda, A = A, lam0 = lam0, lamj = lamj, p0 = p0)R> MTD

### v2

@@ 1 Introduction
Categorical time series with long-range dependencies are ubiquitous in nature. Today’s weather depends not only on the weather of the previous day, but also on what the weather was like a year ago. In large language models, incorporating relevant words from distant positions in the input sequence can substantially enhance next-word prediction (more than $10^{5}$ for ChatGPT 4o). Economic indicators’ dynamics depends on events that happen at different temporal scales, including long-range dependencies. Modeling these phenomena using stochastic processes with possibly long-range dependencies is natural. Given the complex nature of some of these phenomena, often non-parametric modeling is desirable. Nevertheless, non-parametric modeling, like high-order Markov chains, generally requires estimating joint probability distributions, which imposes significant constraints on the size of the past that can be included in the model before the estimation becomes inaccurate. To ameliorate the problem, Raftery85 introduced the mixture transition distribution (MTD) model, a subclass of finite-order Markov chains, as a parsimonious non-parametric model for categorical time series. The MTD framework successfully reduces the number of parameters required for estimation, alleviating the complexity associated with high-order dependencies. Despite this advancement, accurately estimating time series dependencies of an MTD model that stretch far into the past remained limited. Recently, Ost&Takahashi introduced a computationally and statistically efficient solution to the problem of estimating time-series long-range dependencies if the time series is a realization of an MTD model in which the number of relevant lags in the past is not too big. This opened the possibility to make non-parametric inferences for the class of high-order MTD models. In this article, we introduce hdMTD, an R package allowing us to estimate parameters of such high-dimensional Markovian models.
@@
The core of the hdMTD package lies in the algorithms that, given a sample from an MTD chain, can retrieve the relevant set of past lags using two different methods: (1) the Bayesian information criterion (BIC) algorithm, which is a classical model selection method; (2) the forward stepwise and cut (FSC) algorithm, which is a computationally and statistically efficient algorithm for estimating the set of relevant lags even when the temporal distance between the lags is close to the sample size. Once the relevant lags are specified, hdMTD package can also compute the maximum likelihood estimate (MLE) of the transition probabilities and estimate high-order MTD parameters using the expectation-maximization (EM) algorithm. To allow for Monte Carlo experiments, it is necessary to sample from the invariant distribution of the process. For low-order Markov chains, traditionally we rely on running the process with fixed pasts and wait until the process approximate the invariant measure. When the order of the chain is large, this approach can be inefficient as it is difficult to decide whether the joint probability measure is close the to the invariant distribution. A more robust approach is to sample exactly from the invariant distribution. Algorithms that allow to draw samples from the invariant distribution of a process are called perfect sampling algorithms (see (Fernandez2001Coupling) for a comprehensive introduction on the subject). Our package allows to simulate MTD chains of any order from its invariant distribution using the perfect sampling algorithm, which is novel to the best of our knowledge .
@@
A brief review of related computational packages helps to place hdMTD in context. Markovian models have a long history and broad applications, nevertheless, practical tools for inference, particularly for mixture transition distribution (MTD) models, remain relatively limited. Traditional statistical software (e.g., SPSS (SPSS), Stata (Stata)) do not provide native, dedicated support for Markov models, while SAS (SAS) offers limited functionality, primarily for homogeneous chains. Consequently, researchers increasingly rely on open-source environments like R (R-base) for advanced Markovian analysis. In R, the march (march) (Berchtold2020) package provides dedicated tools for MTD model estimation, offering likelihood based optimization for categorical stochastic processes. This package has become a reference for MTD applications, supporting homogeneous Markov chains, hidden Markov models, and other related architectures. Its companion package GenMarkov (GenMarkov) extends this to multivariate MTDs. Notably, none of these tools provide perfect simulation algorithms of MTD processes given parameters. For general Markov chain analysis, the markovchain (markovchain17) package analyzes properties (e.g., stationarity) and experimentally fits $k$-step Markov chains, while DTMCPack (DTMCPack) focuses on simulation from user-defined matrices. The depmixS4 (depmixS4) package offers tools for fitting standard and hidden Markov models with both continuous and categorical responses. Outside R, specialized tools like Stata’s markov module (Statamarkov) implement basic discrete-time Markov chain analysis. More recently, developments in Julia (Julia) and Python (Python) have started to offer basic Markovian model support, but tools specific for MTDs remain largely unavailable.
@@ 2.3 Mixture transition distribution (MTD) models
An MTD (Raftery85; Berchtold&Raftery) of order $d\in\mathbb{Z}^{+}$ is a Markov chain of order $d$ with state space $\mathcal{A}$, where the transition probabilities can be expressed as
@@
Henceforth, an index $j\in\operatorname{\llbracket}-d,0\operatorname{\rrbracket}$ will be called the j-th lag of the model. Throughout this article, for $j\in\operatorname{\llbracket}-d,-1\operatorname{\rrbracket}$, we denote by $p_{j}=\{p_{j}(a|b):\ a,b\in\mathcal{A}\}$, the $|\mathcal{A}|\times|\mathcal{A}|$ stochastic matrix representing the relationship between the $j$-th lag and the present. Equation (2) represents what Berchtold&Raftery called a multimatrix MTD model, since the stochastic matrices $p_{j}$ are allowed to be different from each other. Note that a multimatrix MTD model may have up to $d+(|\mathcal{A}|-1)(1+d|\mathcal{A}|)$ parameters. Hence, for such Markovian models the number of parameters grows linearly with the order $d$, unlike general Markov chains for which the number of parameters grows exponentially with the order. Note also that there are even more parsimonious MTD models, such as the single matrix MTD, where all lags influence the distribution of the present state equally (i.e., given any $a,b\in\mathcal{A}$, $p_{j}(a|b)=p(a|b)$ for all $j\in\operatorname{\llbracket}-d,-1\operatorname{\rrbracket}$). This latter MTD model has only $d+(|\mathcal{A}|-1)(1+|\mathcal{A}|)$ parameters.
@@ 3.2 Empirical probabilities
One can show that, when $\Lambda\subseteq S$, the empirical conditional probability $\hat{\operatorname{\mathsf{P}}}_{n}(a|x_{S})$ defined in (7) converges in probability to the conditional probability $\operatorname{\mathsf{P}}(a|x_{\Lambda})$ as the sample size $n\to\infty$, provided that any two vectors $x,y\in\mathcal{A}^{d}$ of $d$ past states communicate, meaning that the MTD model generating the sample is irreducible (see, e.g., Finesso). Under this irreducibility assumption and if $\hat{\Lambda}_{n}=\hat{\Lambda}_{n}(\mathcal{S}_{n})$ is a consistent estimator of $\Lambda$ (i.e., the probability of the event $\{\hat{\Lambda}_{n}\neq\Lambda\}$ vanishes as the sample size $n\to\infty$), then we also have
@@ 3.3 BIC estimator
Model selection through the BIC (Schwarz) is a well-known method. The difference between BIC values of two models approximates twice the logarithm of the Bayes factor between them. The Bayes factor represents the evidence, provided by the data, in favor of one model against another (Kass&Raftery) and is known for its validity for comparison of multiple and non-nested models. Also, the BIC has been proven to be a consistent method for estimating the order of Markov chains (Katz; Csiszar&Shields).
@@ 3.4 CUT estimator
Ost&Takahashi proposed a consistent estimator for the set of relevant lags of an MTD model based on pairwise comparisons of empirical conditional probabilities corresponding to compatible pasts. In this article, we call it CUT estimator [^2] In Ost&Takahashi, this estimator based on pairwise comparisons is called PCP estimator..
Given a subset $S\subseteq\operatorname{\llbracket}-d,-1\operatorname{\rrbracket}$ and a lag $j\in S$, two sequences $x_{S},y_{S}\in\mathcal{A}^{S}$ are called $(S\setminus\{j\})$-compatible if $x_{k}=y_{k}$ for all $k\in S\setminus\{j\}$. From our discussion at the end of Section 2.3, if $j\in\Lambda\subseteq S$ there must exist a pair of $(S\setminus\{j\})$-compatible pasts $x_{S},y_{S}\in\mathcal{A}^{S}$ such that the total variation distance between $\operatorname{\mathsf{P}}(\cdot|x_{S})$ and $\operatorname{\mathsf{P}}(\cdot|y_{S})$ is strictly positive. Hence, the CUT estimator runs across all $j\in S$, and tests if there is a pair of $(S\setminus\{j\})$-compatible pasts $x_{S},y_{S}\in\mathcal{A}^{S}$ for which
@@
with $\xi>0$, $\alpha>0$ and $\mu\in(0,3)$ such that $\mu>\psi(\mu)=e^{\mu}-\mu-1$. The threshold $t_{n}(x_{S},y_{S})$ has been derived through martingale concentration inequalities (see Ost&Takahashi). An interesting feature of this threshold is that it is adapted to each realization of sample $\mathcal{S}_{n}$ of the MTD model, i.e., it may change depending on the realization of the sample. Using uniform thresholds that do not depend on the sample typically lead to statistical procedures that either underestimate or overestimate the set of relevant lags.
The CUT estimator is computed in the function hdMTD_CUT. In Ost&Takahashi, a few theoretical guarantees of the CUT estimator have been established. Let us highlight two of them. First, the CUT estimator can be implemented with $O(|\mathcal{A}|^{2}|S|(n-d))$ operations (see Ost&Takahashi). Second, in Ost&Takahashi, the CUT estimator is shown to be consistent even when the order $d$ is large (e.g. $d=n\beta$ for some $\beta\in(0,1)$), provided that $\Lambda\subseteq S$ and $|S|$ is relatively small with respect to the sample size $n$ (i.e., $|S|=O(\log(n))$). However, this assumption is too restrictive and might be difficult to be justified in practice. To circumvent this issue, one can use the FS or FSC estimators described subsequently.
@@ 3.5 FS and FSC estimators
The FS estimator (Ost&Takahashi) is defined iteratively as follows. At each step, it receives a subset $S\subseteq\operatorname{\llbracket}-d,-1\operatorname{\rrbracket}$ as input and computes, based on the sample $\mathcal{S}_{n}$, the empirical estimate $\hat{\nu}_{n,j,S}$ of the quantity $\bar{\nu}_{j,S}=\mathbb{E}(\nu_{j,S}(X_{S})),$ for each lag $j\in S^{c}$. Here, for each $x_{S}\in\mathcal{A}^{S}$,
@@
As indicated in Ost&Takahashi, one can compute the FS estimator with $O(|A|^{3}\ell(n-d)(d-(\ell-1)/2)$ operations. Moreover, Theorem 2 of Ost&Takahashi states that, on a certain event and for a suitable chosen $\ell$, we have that $\Lambda\subseteq\hat{\Lambda}_{n}^{\rm FS}$. In hdMTD package, $\ell$ is a user input, and in practice, we verify that if $\ell\geq|\Lambda|$, then $\Lambda\subseteq\hat{\Lambda}_{n}^{\rm FS}$ with high probability. This suggests that the FS output might be a reasonable choice for input set $S$ of both BIC and CUT estimators. Let us mention that using the FS output as the CUT estimator input constitutes what Ost&Takahashi refer to as the FSC estimator. Roughly speaking, the FSC estimator is proved to be consistent even when the underlying MTD model has a large order $d$, as long as the size of the set of relevant lags $\Lambda$ grows at most logarithmically with the sample size $n$. As opposed to the CUT estimator, the FSC estimator does not require any prior knowledge of a sufficiently small subset $S$ containing the set of relevant lags $\Lambda$, which is a huge advantage in practice.
@@ 4 Algorithm to exactly simulate MTD models
There is no exact simulation algorithm implemented in any of the packages for MTD models that we are aware of, namely, the packages march and GenMarkov. In the hdMTD, the above exact simulation algorithm is implemented in the function perfectSample. Our algorithm is inspired by (Comets2002Proceses) - where the authors present an algorithm to exactly simulate processes with long memory. In that paper, the authors also discuss the differences and similarities of their algorithm with respect to other classical exact simulation algorithms such as coupling from the past algorithm introduced by (Propp1996Exact). We also refer to (Fernandez2001Coupling) for a comprehensive introduction to perfect simulation algorithms.
@@ 5 Using hdMTD
R> install.packages("hdMTD")
and loaded with:
@@ 5.1 Data generation
Suppose we want to obtain a sample from an MTD model with state space $\mathcal{A}=\{0,1\}$ where the set of relevant lags is $\Lambda=\{-30,-15,-1\}$. First, we need to specify such a model. We do so with the MTDmodel function. This function creates a class MTD object whose attributes are all the necessary parameters for defining an MTD model, and the resulting transition matrix. The only inputs that are absolutely necessary are the set of relevant lags (Lambda) and the state space (A). Note that while $\Lambda\subset\mathbb{Z}^{-}$, the argument Lambda of the function MTDmodel must be in $\mathbb{Z}^{+}$. Hence, the input set Lambda has to be interperted as the set $\Lambda$. The user can also input the MTD weights (a number lam0 representing $\lambda_{0}$ and a vector lamj representing the values $\lambda_{j}$ for all $j\in\Lambda$), the past conditional distributions (a list pj where each entry is a matrix representing the matrices $p_{j}$, $j\in\Lambda$), and/or the independent distribution (a vector p0, representing $p_{0}(a)$, $a\in\mathcal{A}$). If not provided, these parameters are sampled uniformly and normalized [^3] For sampling $p_{0}(a)$, $a\in\mathcal{A}$, the function first sets p0<-runif(length(A)), and then renormalizes the vector $p_{0}$ to obtain a probability measure by defining p0<-p0/sum(p0). The same procedure is done for the weights (lam0 and lamj). For sampling $p_{j}(a|b)$, $a,b\in\mathcal{A}$, $j\in\Lambda$, the function sets a list called pj with $|\Lambda|$ matrices, where each row of each matrix is sampled in the same way as p0. .
R> set.seed(11)R> Lambda <- c(1, 15, 30)R> A <- c(0, 1)R> lam0 <- 0.01R> lamj <- c(0.39, 0.3, 0.3)R> p0 <- c(0.5, 0.5)R> MTD <- MTDmodel(Lambda = Lambda, A = A, lam0 = lam0, lamj = lamj, p0 = p0)R> summary(MTD)
Mixture Transition Distribution (MTD) modelCall:MTDmodel(Lambda = Lambda, A = A, lam0 = lam0, lamj = lamj, p0 = p0)Relevant lags: -1, -15, -30State space: 0, 1lambdas (weights): lam0 lam-1 lam-15 lam-30 0.01 0.39 0.30 0.30Independent distribution p0:p0(0) p0(1) 0.5 0.5Transition matrices pj (one per lag): pj for lag j = -1: 0 10 0.35190318 0.64809681 0.03558321 0.9644168 pj for lag j = -15: 0 10 0.4278830 0.57211701 0.7670555 0.2329445 pj for lag j = -30: 0 10 0.8341439 0.16585611 0.2184814 0.7815186Transition matrix P: 8 x 2- Preview of first rows of P: 0 1000 0.5208503 0.4791497001 0.3974855 0.6025145010 0.6226020 0.3773980011 0.4992372 0.5007628100 0.3361516 0.6638484101 0.2127868 0.7872132Reading guide for P:Rows list past contexts from oldest to newest, matching lags (-30, -15, -1).
@@
In addition to the summary() method, the MTD object includes a set of accessor functions and methods that facilitate model exploration and manipulation. Accessor functions such as transitP(), lambdas(), pj(), p0(), lags(), Lambda(), and states() return the core components of the model in a structured way. The object also supports various S3 methods, including coef(), logLik(), probs(), perfectSample(), and oscillation()—the last two being further discussed later in this section. To guide the user, the print() method displays a concise overview of the model structure, along with the full list of available accessors and methods.